# Weingarten calculus with virtual isometries

### v2

@@ 2.5. Calculation of Haar-distributed matrices from reflection matrices
Let $X_{1},X_{2},\dots,X_{n}$ be independent random vectors, where each $X_{k}$ is uniformly distributed on the unit sphere of $\mathbb{C}^{k}$. For each $1\leq k\leq n$, we denote by $R_{k}$ the corresponding reflection matrix of size $k$ given in a manner similar to (2.1), and we define an $n\times n$ matrix $\tilde{R}_{k}=R_{k}\oplus I_{n-k}$. Then, a Haar-distributed unitary matrix $U$ on $\mathrm{U}(n)$ has the same distribution as the product $\tilde{R}_{n}\tilde{R}_{n-1}\cdots\tilde{R}_{1}$. In particular, the last $k$ columns of $U$ match those of $\tilde{R}_{n}\tilde{R}_{n-1}\cdots\tilde{R}_{n-k+1}$.
Thanks to this lemma, it is possible to calculate moments of polynomial functions on the unitary group by a method different from the Weingarten calculus, using Theorem 2.4. In this subsection, we provide some examples of such calculations, that indicate that the approach described in this paper allows to recover Weingarten functions.
@@
Using the Weingaten calculus (1.1), it is immediately clear that this value is equal to $\operatorname{Wg}_{2,n}((1\ 2))=-\frac{1}{n(n+1)(n-1)}$, but here we shall deliberately take a longer route to calculate it. Since $(r_{ij})$ and $(s_{ij})$ are independent, we have
@@
Example 2.11 or Proposition 2.7 imply that
@@
Using Proposition 2.10 (see also Example 2.11), we have
$\mathbb{E}[r_{n-1,j}r_{n,n}\overline{r_{n-1,j}r_{n,n}}]=\begin{cases}\frac{1}{n^{2}(n+1)}&\text{if $j<n-1$};\\ \frac{n^{2}-n+2}{n^{2}(n+1)}&\text{if $j=n-1$},\end{cases}$
which gives
@@
This approach works for any integral of a monomial in $u_{ij}$ and their conjugates that has a single matching, resulting in a Weingarten function, and this provides a method to recompute the Weingarten function. It remains to identify the underpinning algebraic and combinatorial phenomena that allow one to rederive systematically the known results on Weingarten functions, which we leave as an open problem. For a related result, we also refer to Proposition 3.11.
@@ 3.1. A recursive formula
and it follows that $f_{1}*f_{2}=f_{2}*f_{1}$. We use such substitutions without notice.
@@
Recall the Weingarten function already appeared in the introduction. Suppose that $k\leq n$. The Weingarten function for the unitary group $\mathrm{U}(n)$ is given by
@@
for each permutation $\pi\in S_{k}$, where $U=(u_{ij})$ is a Haar-distributed unitary matrix in $\mathrm{U}(n)$. The origin of the name of $\operatorname{as}_{k,n}$ is based on the following theorem.
@@
Theorem 3.3 does not hold without the assumption $k+1\leq n$. In the case where $k+1>n$, the Weingarten function $\operatorname{Wg}_{k,n-1}$ on the right-hand side is not well defined. Although the Weingarten functions can still be extended in that case, the above formula does not hold. We will elaborate on these points in Section $3.$
@@
Expanding the factors, we have
@@
where $\operatorname{Fix}(\sigma)^{c}:=\{1,2,\dots,k\}\setminus\operatorname{Fix}(\sigma)$ is invariant under $\sigma$. By Lemma 2.2, each factor in the last sum is equal to $1/n^{\uparrow(s+k-\mathsf{f}(\sigma))}$. Therefore, we have
@@
Since we are currently assuming $k+1\leq n$, the elements of $U$ in the $n$-th row or $n$-th column do not appear in this equation. By the independence of $R$ and $V$, we have
@@
Next, we apply the Weingarten calculus (1.1) for $\mathrm{U}(n-1)$ to the last factor:
@@
or $\operatorname{Wg}_{k,n}(\pi)=(\widetilde{\operatorname{as}}*\operatorname{Wg}_{k,n-1})(\pi)$ with
@@
The final step in the proof is to prove that $\widetilde{\operatorname{as}}(\sigma)$ is equal to $\operatorname{as}_{k,n}(\sigma)$, as defined in Definition 3.1. We start with the expression for $\widetilde{\operatorname{as}}(\sigma)$:
@@
Since $R=(r_{ij})$ is an $n\times n$ unitary matrix, we can analyze the sums involved. Specifically, we have:
@@
for all $1\leq i\leq k$. From (2.1), we know that $r_{i,n}=x_{i}$. Substituting this into the above equation, we obtain:
@@
This equation, combined with Lemma 3.5, shows that $\widetilde{\operatorname{as}}(\sigma)=\operatorname{as}_{k,n}(\sigma)$. This completes the proof of Theorem 3.3. ∎
@@ 3.2. Some properties for ascension functions
In this subsection, we investigate some properties of the function $\operatorname{as}_{k,n}:S_{k}\to\mathbb{C}$. To do this, we review some fundamental properties of the Weingarten function, refer to [CMN22] and its references for details.
@@
where $\kappa(\pi)$ is the number of cycles in the cycle decomposition of $\pi$. This function is invertible if and only if $k\leq n$, and in this case, the inverse function of $\operatorname{G}_{k,n}$ is the Weingarten function $\operatorname{Wg}_{k,n}$:
@@
Here $\delta_{e_{k}}:S_{k}\to\{0,1\}$ is the Dirac function at the identity permutation $e_{k}$:
@@
From Theorem 3.3, it is immediate to see the formula
@@
Note that we need to assume $k\leq n$ in the latter equation to ensure that the factor $(n\uparrow\lambda)$ of all terms does not equal 0. From these formulas, we can easily obtain the character expansion of the ascension function using the orthogonality of irreducible characters
@@
(3.3) $\operatorname{as}_{k,n}=\frac{1}{k!}\sum_{\lambda\vdash k}f^{\lambda}\left(\prod_{i=1}^{d(\lambda)}\frac{n-1-(\lambda_{i}^{\prime}-i)}{n-(\lambda_{i}-i)}\right)\chi^{\lambda},$
where $(\lambda^{\prime}_{1},\lambda_{2}^{\prime},\dots)$ represents the conjugate partition of $\lambda$, obtained by transposing the partition when viewed as a Young diagram along its diagonal. Additionally, $d(\lambda)=|\{i\geq 1\ |\ \lambda_{i}\geq i\}|$ denotes the length of the diagonal in the Young diagram $\lambda$. Here, the ratio in the coefficient can be derived from the following identity:
$\frac{((n-1)\uparrow\lambda)}{(n\uparrow\lambda)}=\prod_{i=1}^{d(\lambda)}\frac{n-1-(\lambda_{i}^{\prime}-i)}{n-(\lambda_{i}-i)},$
@@
Let $k$ and $n$ be positive integers. We assume $k\leq n$ and define the descension function $\operatorname{ds}_{k,n}$ in $\mathbb{C}[S_{k}]$ by
@@
The falling factorial term imposes the assumption $k\leq n$.
@@
Let $k$ and $n$ be positive integers, and we assume $k\leq n$. Then it holds that
@@
The second formula follows from Theorem 3.3 and (3.4). To show the first formula (3.4), we extend the range of the parameter of the functions $\operatorname{G}_{k,n}$, $\operatorname{Wg}_{k,n}$, and $\operatorname{as}_{k,n}$ from a positive integer $n$ (where $k\leq n$) to a complex number $z$, as referenced in [CŚ06]. We define $\operatorname{G}_{k,z}$ by
@@
This function has a trivial symmetry:
$\operatorname{G}_{k,-z}(\sigma)=(-1)^{\kappa(\sigma)}\operatorname{G}_{k,z}(\sigma)=(-1)^{k}\operatorname{sgn}(\sigma)\operatorname{G}_{k,z}(\sigma)$
since $\operatorname{sgn}\sigma=(-1)^{k-\kappa(\sigma)}$. It is known that $\operatorname{G}_{k,z}$ is invertible in the group algebra $\mathbb{C}[S_{k}]$ if and only if $z\not\in\{0,1,\dots,k-1\}$. This fact can be confirmed from (3.2). In this case, we denote by $\operatorname{Wg}_{k,z}$ the inverse of $\operatorname{G}_{k,z}$. It also satisfies the symmetry relation $\operatorname{Wg}_{k,-z}(\sigma)=(-1)^{k}\operatorname{sgn}(\sigma)\operatorname{Wg}_{k,z}(\sigma)$.
@@
based on (3.1). This function is invertible if and only if $z\not\in\{1,\dots,k\}$. In particular, if $k\leq n$ then $\operatorname{as}_{k,n+1}$ is invertible. Since $\operatorname{G}_{k,z}$ and $\operatorname{Wg}_{k,z}$ are inverses of each other for all $z\not\in\{0,1,\dots,k-1\}$, when we denote the inverse of $\operatorname{as}_{k,z+1}$ as $\operatorname{as}^{-1}_{k,z+1}$, we can transform it as follows using symmetries:
@@
Therefore, the inverse of $\operatorname{as}_{k,n+1}$ is $\operatorname{sgn}*\operatorname{as}_{k,-n}$, which is equal to $\operatorname{ds}_{k,n}$. Indeed, we have:
@@
We have proved Theorem 3.3 using random matrices; however, since it is an algebraic equation, it would be interesting to look for an algebraic proof.
@@ 3.3. Weingarten functions in the case of low dimensions
In the previous discussions, we often assumed that $k\leq n$. This is due to the fact that the function $\operatorname{G}_{k,n}(\sigma)=n^{\kappa(\sigma)}$ on $S_{k}$ is invertible only when $k\leq n$, and the inverse element is $\operatorname{Wg}_{k,n}$. The Weingarten calculus itself (1.1) holds even when $k>n$; however, in that case, it is necessary to replace $\operatorname{Wg}_{k,n}$ with the next $w_{k,n}$. The function $w_{k,n}$ on $S_{k}$ is a class function satisfying the following equation:
@@
Such functions $w_{k,n}$ are not uniquely defined if $k>n$, but they are also referred to as Weingarten functions. When $k\leq n$, $w_{k,n}$ is uniquely determined and coincides with $\operatorname{Wg}_{k,n}$.
@@
Assume that $k>n$. If we also assume $w_{k,n}*\operatorname{G}_{k,n}*w_{k,n}=w_{k,n}$ in addition to (3.5), then $w_{k,n}$ is uniquely determined. (This condition was insufficient in [Mat13].) If we denote such a $w_{k,n}$ specifically as $W_{k,n}$, then $W_{k,n}$ is concretely given by
@@
which includes a slight restriction $\ell(\lambda)\leq n$ added to (3.2). Here, $\ell(\lambda)=|\{i\geq 1\ |\ \lambda_{i}>0\}|$ is the length of a partition $\lambda=(\lambda_{1},\lambda_{2},\dots)$. As already mentioned, the Weingarten formula (1.1) holds for any $w_{k,n}$ that satisfies (3.5), not just for this special $W_{k,n}$ – we refer to [CFM25] for further details.
Unfortunately, Theorem 3.3 does not hold for functions $w_{k,n}$ when $k\geq n$. Indeed, let $k=n=2$ and $\sigma=e_{2}\in S_{2}$. If we apply Theorem 3.3 formally in this case, it is expected that the following holds:
@@
It is easy to see that $w_{2,2}(e_{2})=\operatorname{Wg}_{2,2}(e_{2})=\frac{1}{3}$, while Definition 3.1 and (3.5) imply $\sum_{\pi\in S_{2}}\operatorname{as}_{2,2}(\pi)w_{2,1}(\pi)=\frac{1}{6}\sum_{\pi\in S_{2}}w_{2,1}(\pi)=\frac{1}{12}.$ Therefore, for the case when $k\geq n$, a different equation corresponding to Theorem 3.3 needs to be sought. In the general case, it is not possible, but for Theorem 3.3 when $k=n$, we could obtain a similar expression as follows.
@@
(3.7) $a_{n}(\pi,\tau)=\sum_{t=0}^{\mathsf{f}(\pi,\tau)}(-1)^{n-t+1}\binom{\mathsf{f}(\pi,\tau)}{t}\frac{1}{n^{\uparrow(n-t)}}$
@@
In this theorem, it should be noted that $a_{n}(\pi,\tau)$ is slightly different from $\operatorname{as}_{n,n}(\pi\tau)$.
@@
both of which are equal to $\frac{1}{3}$. As another example of the theorem, the following equation holds true:
@@
When substituting specific values, this equation becomes:
@@
The proof is almost the same as that of Theorem 3.3. we will focus on explaining the different parts. Let the matrices $U=(u_{ij})$, $R=(r_{ij})$, and $V=(v_{ij})$ be as stated in Lemma  2.1. Let $\pi\in S_{n}$, and we consider
@@
$u_{ij}=\begin{cases}\sum_{p=1}^{n-1}r_{i,p}v_{p,j}&\text{if $1\leq i\leq n$ and $1\leq j\leq n-1$};\\ r_{i,n}&\text{if $1\leq i\leq n$ and $j=n$}.\end{cases}$
@@
$F=\{i\in\{1,2,\dots,n-1\}\ |\ \pi(\tau(i))=i\}|,$
and put $F^{\prime}=\{1,2,\dots,n\}\setminus F$. Note that $F^{\prime}$ is not the complement in $\{1,2,\dots,n-1\}$. Then, we have:
@@
By combining these, the Weingarten function $\operatorname{Wg}_{k,n}\in\mathbb{C}[S_{k}]$ with $k\leq n$ can be expressed using two types of functions: $\operatorname{as}_{k,*}(\sigma)$ and $a_{*}(\sigma,\tau)$.
@@
This theorem allows the Weingarten function to be calculated sequentially starting from $n=1$.

### v3

@@ 2.5. Calculation of Haar-distributed matrices from reflection matrices
Let $X_{1},X_{2},\dots,X_{n}$ be independent random vectors, where each $X_{k}$ is uniformly distributed on the unit sphere of $\mathbb{C}^{k}$. For each $1\leq k\leq n$, we denote $R_{k}$ as the corresponding reflection matrix of size $k$, defined similarly to (2.1), and define an $n\times n$ matrix $\tilde{R}_{k}=R_{k}\oplus I_{n-k}$. Then, a Haar-distributed unitary matrix $U$ in $\mathrm{U}(n)$ has the same distribution as the product $\tilde{R}_{n}\tilde{R}_{n-1}\cdots\tilde{R}_{1}$. In particular, the last $k$ columns of $U$ match those of $\tilde{R}_{n}\tilde{R}_{n-1}\cdots\tilde{R}_{n-k+1}$.
Using this lemma, one can calculate moments of polynomial functions on the unitary group $\mathrm{U}(n)$ by a method different from the Weingarten calculus, employing Theorem 2.4. In this subsection, we provide some examples of such calculations, which indicate that the approach described in this paper allows one to recover Weingarten functions.
@@
By applying the Weingarten calculus (1.1), one immediately obtains
$\mathbb{E}\left[u_{n-1,n-1}u_{n,n}\overline{u_{n-1,n}u_{n,n-1}}\right]=\operatorname{Wg}_{2,n}((1\ 2))=-\frac{1}{n(n+1)(n-1)},$
but here, we shall deliberately take a longer route to calculate it. Noting that $(r_{ij})$ and $(s_{ij})$ are independent, it follows that
@@
Example 2.11 or Proposition 2.7 give that
@@
Using Proposition 2.10 (see also Example 2.11), we have:
$\mathbb{E}[r_{n-1,j}r_{n,n}\overline{r_{n-1,j}r_{n,n}}]=\begin{cases}\frac{1}{n^{2}(n+1)}&\text{if $j<n-1$};\\ \frac{n^{2}-n+2}{n^{2}(n+1)}&\text{if $j=n-1$}.\end{cases}$
This gives
@@
This approach works for any integral of a monomial in $u_{ij}$ and their conjugates, provided it has a single matching. Such an integral results in a Weingarten function, and this provides a method to recompute the Weingarten function
It remains to identify the underpinning algebraic and combinatorial phenomena that allow one to systematically rederive the known results on Weingarten functions, which we leave as an open problem. For a related result, we also refer to Proposition 3.11.
@@ 3.1. A recursive formula
and it follows that $f_{1}*f_{2}=f_{2}*f_{1}$. We will use this commutativity property without further notice.
@@
Recall the Weingarten function which already appeared in the introduction. Suppose that $k\leq n$. The Weingarten function for the unitary group $\mathrm{U}(n)$ is given by
@@
for each permutation $\pi\in S_{k}$, where $U=(u_{ij})$ is a Haar-distributed unitary matrix in $\mathrm{U}(n)$. The name ‘ascent function’ for $\operatorname{as}_{k,n}$ originates from the following theorem.
@@
Theorem 3.3 does not hold without the assumption $k+1\leq n$. In the case where $k+1>n$, the Weingarten function $\operatorname{Wg}_{k,n-1}$ on the right-hand side is not well-defined. Although the Weingarten functions can still be extended in that case, the above formula does not hold. We will elaborate on these points in Section $3.$
@@
By expanding the product, we have
@@
where $\operatorname{Fix}(\sigma)^{c}:=\{1,2,\dots,k\}\setminus\operatorname{Fix}(\sigma)$ is invariant under $\sigma$. By Lemma 2.2, each expectation appearing in the sum equals $1/n^{\uparrow(s+k-\mathsf{f}(\sigma))}$. Therefore, we obtain
@@
Since we assume $k+1\leq n$, the elements of $U$ in the $n$-th row or $n$-th column do not appear in this equation. By the independence of $R$ and $V$, we have
@@
Next, we apply the Weingarten formula (1.1) for the unitary group $\mathrm{U}(n-1)$ to the last factor:
@@
which can be written as $\operatorname{Wg}_{k,n}(\pi)=(\widetilde{\operatorname{as}}*\operatorname{Wg}_{k,n-1})(\pi)$ with
@@
The final step in the proof is to show that $\widetilde{\operatorname{as}}(\sigma)$ equals $\operatorname{as}_{k,n}(\sigma)$, as defined in Definition 3.1. Starting from the expression for $\widetilde{\operatorname{as}}(\sigma)$, we have
@@
Since $R=(r_{ij})$ is an $n\times n$ unitary matrix, the sums can be simplified as follows:
@@
for all $1\leq i\leq k$. From (2.1), we know that $r_{i,n}=x_{i}$. Substituting this, we obtain
@@
Together with Lemma 3.5, this shows $\widetilde{\operatorname{as}}(\sigma)=\operatorname{as}_{k,n}(\sigma)$. This completes the proof of Theorem 3.3. ∎
@@ 3.2. Some properties for ascension functions
In this subsection, we investigate some properties of the function $\operatorname{as}_{k,n}:S_{k}\to\mathbb{C}$. To do this, we review some fundamental properties of the Weingarten function; refer to [CMN22] and its references for details.
@@
where $\kappa(\pi)$ denotes the number of cycles in the cycle decomposition of $\pi$. This function is invertible if and only if $k\leq n$, and in this case, its inverse function is the Weingarten function $\operatorname{Wg}_{k,n}$:
@@
Here, $\delta_{e_{k}}:S_{k}\to\{0,1\}$ is the Dirac function at the identity permutation $e_{k}$:
@@
From Theorem 3.3, it immediately follows that
@@
Note that we need to assume $k\leq n$ in the latter equation to ensure that the factor $(n\uparrow\lambda)$ in all terms is not equal to zero. From these formulas, we can easily obtain the character expansion of the ascension function by using the orthogonality of irreducible characters:
@@
(3.3) $\operatorname{as}_{k,n}=\frac{1}{k!}\sum_{\lambda\vdash k}f^{\lambda}\left(\prod_{i=1}^{d(\lambda)}\frac{n-1-(\lambda_{i}^{\prime}-i)}{n+(\lambda_{i}-i)}\right)\chi^{\lambda},$
where $(\lambda^{\prime}_{1},\lambda_{2}^{\prime},\dots)$ represents the conjugate partition of $\lambda$, obtained by transposing the partition, viewed as a Young diagram, along its diagonal. Additionally, $d(\lambda)=|\{i\geq 1\ |\ \lambda_{i}\geq i\}|$ denotes the length of the diagonal in the Young diagram $\lambda$. Here, the ratio in the coefficient can be derived from the following identity:
$\frac{((n-1)\uparrow\lambda)}{(n\uparrow\lambda)}=\prod_{i=1}^{d(\lambda)}\frac{n-1-(\lambda_{i}^{\prime}-i)}{n+(\lambda_{i}-i)},$
@@
Let $k$ and $n$ be positive integers such that $k\leq n$. We define the descension function $\operatorname{ds}_{k,n}$ in $\mathbb{C}[S_{k}]$ by
@@
Note that the definition of the falling factorial requires the assumption $k\leq n$.
@@
Let $k$ and $n$ be positive integers such that $k\leq n$. Then the following holds:
@@
The second formula follows from Theorem 3.3 and (3.4). To show the first formula (3.4), we extend the parameter range of the functions $\operatorname{G}_{k,n}$, $\operatorname{Wg}_{k,n}$, and $\operatorname{as}_{k,n}$ from a positive integer $n$ (with $k\leq n$) to a complex number $z$, following [CŚ06]. We define $\operatorname{G}_{k,z}$ by
@@
This function satisfies the following symmetry:
$\operatorname{G}_{k,-z}(\sigma)=(-1)^{\kappa(\sigma)}\operatorname{G}_{k,z}(\sigma)=(-1)^{k}\operatorname{sgn}(\sigma)\operatorname{G}_{k,z}(\sigma),$
since $\operatorname{sgn}\sigma=(-1)^{k-\kappa(\sigma)}$. It is known that $\operatorname{G}_{k,z}$ is invertible in the group algebra $\mathbb{C}[S_{k}]$ if and only if $z\not\in\{0,1,\dots,k-1\}$. This fact can be confirmed by (3.2). In this case, we denote by $\operatorname{Wg}_{k,z}$ the inverse of $\operatorname{G}_{k,z}$. It also satisfies the symmetry relation $\operatorname{Wg}_{k,-z}(\sigma)=(-1)^{k}\operatorname{sgn}(\sigma)\operatorname{Wg}_{k,z}(\sigma)$.
@@
based on (3.1). This function is invertible if and only if $z\not\in\{1,\dots,k\}$. In particular, if $k\leq n$, then $\operatorname{as}_{k,n+1}$ is invertible. We denote the inverse of $\operatorname{as}_{k,z+1}$ by $\operatorname{as}^{-1}_{k,z+1}$. Since $\operatorname{G}_{k,z}$ and $\operatorname{Wg}_{k,z}$ are inverses of each other for all $z\not\in\{0,1,\dots,k-1\}$, by using symmetries, we can transform it as follows:
@@
Therefore, the inverse of $\operatorname{as}_{k,n+1}$ is $\operatorname{sgn}*\operatorname{as}_{k,-n}$, which equals $\operatorname{ds}_{k,n}$. Indeed, we have:
@@
We have proved Theorem 3.3 using random matrices; however, since the theorem is an algebraic statement, it would be interesting to look for an algebraic proof.
@@ 3.3. Weingarten functions in the case of low dimensions
In the previous discussions, we often assumed that $k\leq n$. This is because the function $\operatorname{G}_{k,n}(\sigma)=n^{\kappa(\sigma)}$ on $S_{k}$ is invertible only when $k\leq n$, and the inverse element is $\operatorname{Wg}_{k,n}$. The Weingarten calculus itself (1.1) holds even when $k>n$; however, in that case, it is necessary to replace $\operatorname{Wg}_{k,n}$ with a slightly different function $w_{k,n}$. The function $w_{k,n}$ on $S_{k}$ is a class function satisfying the following equation:
@@
Such functions $w_{k,n}$ are not uniquely defined if $k>n$, but they are also called Weingarten functions. When $k\leq n$, $w_{k,n}$ is uniquely determined and coincides with $\operatorname{Wg}_{k,n}$.
@@
Let us assume that $k>n$. In addition to (3.5), if we have $w_{k,n}*\operatorname{G}_{k,n}*w_{k,n}=w_{k,n}$, then $w_{k,n}$ is uniquely determined. (Note that this condition was insufficient in [MAT13].) If we denote such a $w_{k,n}$ specifically as $W_{k,n}$, then $W_{k,n}$ is concretely given by
@@
which imposes a slight restriction $\ell(\lambda)\leq n$ on (3.2). Here, $\ell(\lambda)=|\{i\geq 1\ |\ \lambda_{i}>0\}|$ is the length of a partition $\lambda=(\lambda_{1},\lambda_{2},\dots)$. As already mentioned, the Weingarten formula (1.1) holds for any $w_{k,n}$ satisfying (3.5), not only for this particular $W_{k,n}$. For more details, see [CFM25].
Unfortunately, Theorem 3.3 does not hold for functions $w_{k,n}$ when $k\geq n$. Consider the case $k=n=2$ and $\sigma=e_{2}\in S_{2}$. Formally applying Theorem 3.3 in this case suggests the following:
@@
It follows that $w_{2,2}(e_{2})=\operatorname{Wg}_{2,2}(e_{2})=\frac{1}{3}$, whereas Definition 3.1 and (3.5) imply $\sum_{\pi\in S_{2}}\operatorname{as}_{2,2}(\pi)w_{2,1}(\pi)=\frac{1}{6}\sum_{\pi\in S_{2}}w_{2,1}(\pi)=\frac{1}{12}.$ Therefore, in the case $k\geq n$, a different equation analogues to Theorem 3.3 must be found. In general, it is not possible; however when $k=n$, a similar expression to Theorem 3.3 can be obtained as follows.
@@
(3.7) $a_{n}(\pi,\tau)=\sum_{t=0}^{\mathsf{f}(\pi,\tau)}(-1)^{n-t+1}\binom{\mathsf{f}(\pi,\tau)}{t}\frac{1}{n^{\uparrow(n-t)}},$
@@
Note that $a_{n}(\pi,\tau)$ differs slightly from $\operatorname{as}_{n,n}(\pi\tau)$.
@@
and both values are equal to $\frac{1}{3}$. As another example, the following equation holds:
@@
By substituting specific values, this equation becomes:
@@
The proof is almost the same as that of Theorem 3.3. We will focus on explaining the different parts. Let the matrices $U=(u_{ij})$, $R=(r_{ij})$, and $V=(v_{ij})$ be as stated in Lemma  2.1. Let $\pi\in S_{n}$, and consider
@@
$u_{ij}=\begin{cases}\sum_{p=1}^{n-1}r_{i,p}v_{p,j}&\text{if $1\leq i\leq n$ and $1\leq j\leq n-1$},\\ r_{i,n}&\text{if $1\leq i\leq n$ and $j=n$}.\end{cases}$
@@
$F=\{i\in\{1,2,\dots,n-1\}\mid\pi(\tau(i))=i\},$
and put $F^{\prime}=\{1,2,\dots,n\}\setminus F$. Note that $F^{\prime}$ is not the complement in $\{1,2,\dots,n-1\}$, but rather in $\{1,2,\dots,n\}$. Then, we have:
@@
By combining these results, the Weingarten function $\operatorname{Wg}_{k,n}\in\mathbb{C}[S_{k}]$ for $k\leq n$ can be expressed in terms of two types of functions: $\operatorname{as}_{k,*}(\sigma)$ and $a_{*}(\sigma,\tau)$.
@@
This proposition allows the Weingarten function to be calculated sequentially starting with the base case $n=1$.